40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue May 7 03:14:06 2013, Jenne, Update, SUS, PRM SUS_LSC violin (FM5) set to correct frequency 
    Reply  Wed Aug 21 23:50:40 2013, Koji, Update, SUS, PRM SUS_LSC violin (FM5) set to correct frequency 
       Reply  Fri Aug 23 01:52:34 2013, Jenne, Update, SUS, Violin filters moved to LSC, from SUS MICHloopNotches_23Aug2013.pdf
          Reply  Fri Aug 23 14:19:28 2013, rana, Update, SUS, Violin filters moved to LSC, from SUS 
Message ID: 9058     Entry time: Fri Aug 23 14:19:28 2013     In reply to: 9057
Author: rana 
Type: Update 
Category: SUS 
Subject: Violin filters moved to LSC, from SUS 

  Just to rephrase somewhat:

  • We balance the BS & PRM actuators in the LSC Output matrix so that there is no MICH signal going into REFL_I @ 100 Hz.
  • REFL Phase is adjusted by driving PRM and minimizing Q/I to within ~1 deg (Q/I ratio of ~2%)
  • The REFL_I sensing matrix element is ~50x bigger than REFL_Q (in W/m)
  • When we turn ON the violin mode notch in the BS SUS, it changes the MICH drive into a purely PRCL drive at that frequency !!!
  • So, putting notches into the SUS screens is bad since it imbalances the drives.
ELOG V3.1.3-